ABC stock is priced at $74.20 per share. The company just paid its $1.10 quarterly dividend.

Annual interest rates are 6.0%. A $70.00 strike European call, maturing in 6 months, sells for

$6.50. Calculate the price of a $70.00 strike European put, maturing in 6 months.
Answer #1

we know that put call parity euqation is

'' value of call(Vc)+ present value of strike (X) = value of put (Vp) + value of stock (Vs)''

When there is a dividend, then

''Vc + X = Vp + Vs - P.V (DIVIDEND)''

Value of Put (Vp) = Vc + X - Vs + P.V(dividend)

a a value of put (Vp) - Vc + X - v(s) + p-vad) - 6 so +(10 xe ). 1 - ao t[-10 xe +[110xéo] = 6.50 + [70% -0303] - th 20+ [1ox